OpenStudy (anonymous):

Multiply and simplify: 8/x^2 x^4/4x A)4x^2 B)2x C) x

OpenStudy (anonymous):

\[\LARGE {8\over x^2}\cdot {x^4\over 4x }\] is there multiplication between fractions? (:

OpenStudy (anonymous):

Yesss! Just like how you have it! :)

OpenStudy (anonymous):

ok... first you are supposed to know this formula: \[\LARGE {8\over x^2 }\cdot {x^4\over 4x}={8x\over 4}\cdot {\cancel{x^3}\over \cancel{x^3} } \] what do you get now ? :)

OpenStudy (anonymous):

\[8\over x^2\]* \[ x^4\over4x\]